A depar ment store paid S41.89 or a dinner plata set verhead expense s 19% of the regular selling price and profit s 19% of the regular selling pnce During a clearance sa the set as sold at a markdown of 24% what was the operating profit or loss on the sa The operatingwas Round the final answer to the nearast cent as needed Round all intermediats values to six decimal places as needed)
Answer #1
Cost of goods $41.89
Assuming Regular Sales Price=S
Overhead expense=0.19*S
Profit=0.19*S
Profit=(Sales Price)-(Cost of goods)-(Overhead expense)
Profit=S-41.89-0.19*S
0.19*S=S-41.89-0.19*S
S-0.19*S-0.19*S=41.89
0.62*S=41.89
S=41.89/0.62= $    67.56
Sales Price for clearance sale(24% markdown) $    51.35 (67.56*(1-0.24))
Cost of goods $41.89
Gross Profit $       9.46 (51.35-41.89)
Overhead expense=0.19*S= $    12.84
Operating Profit $    (3.38) (9.46-12.84)
Operating Loss $       3.38